Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

P68-505: The local transient Universe

PI: E. Kankare, M. Stritzinger, T. Reynolds

ToO + Soft ToO triggers

Observations will be requested with ALFOSC and the OB generator will be used to request all ToO and SoftToO observations.

To summarise:

Go to the directory
  cd ~/scripts/68-505
List the available scripts
  ls -lrt

and execute the OB generated scripts requested in the trigger email.


Fixed half-nights service observations with NOTCam

The OB generator will be used to request all observations. For each half night, the PI will send a 'trigger' through the OB-generator. Note that the trigger needs to be Compiled and/or Queued by the staff observer, or the staff person preparing the observations, using the OB-generator.

In order to queue the observations, enter the OB generator and click on the Queue it! button.
Then compile the scripts manually, either through the compile buttons in the OB generator or from within the OB queue.

Scripts are placed in the directory ~/scripts/68-505/ as above for ALFOSC ToO/Soft ToO.

When an OB is queued, an e-mail will be sent automatically to the observer, which contains all the information about the requested observations, the names and locations of the scripts, instructions for observing and comments from the creator of the trigger.


General information on the ALFOSC spectroscopy and photometry:

  • Possible required setup:
    • Filters: ugriz, (UBVRI)
    • Grisms: #4, #7, #8, #18, #19, #20
    • Slits: 0.75", 1.0", 1.3"

  • If the seeing is better than ~1.2", the 1.0" slit should be used, otherwise the 1.3" one.

  • Take 3 imaging flat fields in the required broadband filters using twilight sky, either in the evening or in the morning, using the alfosc.easyflat script.

  • The exposure time of the spectrum depends on the magnitude of the target:
    • If V < 17 mag the on-source exposure time is 1 x 900 sec
    • If V = 17-18 mag the on-source exposure time is 1 x 1800 sec
    • If V = 18-19 mag the on-source exposure time is 1 x 2400 sec
    • If V > 19 mag the on-source exposure time is 1 x 3600 sec

  • Standards will usually be selected by the triggering individual and OB scripts for observing these will also be present in ~/scripts/66-506. Typically one of the following spectroscopic standards is used (sorted in the order of preference):

  • ING ID Name RA (J2000.0) DEC (J2000.0) Exp. time FC
    SP1036+433 Feige 34 10:39:36.7 +43:06:10.1 100 sec Feige34.gif
    SP1234+254 Feige 66 12:37:23.6 +25:04:00.3 50 sec Feige66.gif
    SP0501+527 G191B2B 05:05:30.6 +52:49:56.0 120 sec G191B2B.gif
    SP1550+330 BD+33 2642 15:51:59.9 +32:56:53.0 60 sec BD+33.2642.gif
    SP2148+286 BD+28 4211 21:51:11.1 +28:51:52.0 60 sec BD+28.4211.gif


  • No blocking filter should be used in any of the spectroscopic observations.

  • The ALFOSC specific Cookbook can be accessed here.


General information on the NOTCam photometry:

NOTCam will be used only in the Service part of this proposal. The service time is on fixed nights and the tracking log will not include these observations.

The OB generator will be used to prepare all observations, and the information below is only for orientation.

  • Possible required setup:
    • Filters: JHKs
    • Camera: WF

  • Take 2 sets (bright and faint with the same exposure time) of 8 imaging flat fields in the required broadband filters using twilight sky, either in the evening or in the morning, using the notcam.skyflat script.

  • Please, make sure that the target counts (including sky background) remain linear, i.e. <25000 counts – check & adjust exposure times if necessary. This is particularly important in the Ks-band.

  • Generally the targets are expected to be distant enough that off-source dithering for background subtraction will not be required and that the on-source exposures can be used also as sky frames. By default, the observing scripts are based on the 9point script which carries out a 9 point dither (3x3 grid) observation. However, if separate sky frames are required a beamswitch script will be used.

  • The NOTCam specific Cookbook can be accessed here.


P68 tracking


Program Tracking
Proposal 68-505
Semester P69: 20240401 - 20241001
Principal Investigator Erkki Kankare, Maximilian Stritzinger, Thomas Reynolds
Programme The local transient Universe
Instrument ALFOSC+NOTCam+FIES
Time Allocation 10 hard ToOs 10h, rest soft 60h
Total Nights 7.0
Type ToO+SoftToO
Notes Override. Soft triggers during Nordic service or Technical nights
Total Time Spent 26:30 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
68-505 2024-04-01 00:41 SoftToO 2023ldh ALHd010114 - ALHd010125, ALHd010135 - ALHd010137 Target observed successfully in all filters.OB 20240401_too_2023ldh_phot was executed with clear skies, 20240401_too_2023ldh_phot_z through thin clouds. Both with seeing around 1 arcsec. Observed
68-505 2024-04-11 00:23 SoftToO HD93521, SN2024cld ALHd110124 - ALHd110133, ALHd110206 - ALHd110210 Std star and template_spec_1_3 OB ran for SN2024cld. Targets observed in dusty conditions within seeing constraints of the OBs. Observed
68-505 2024-04-12 00:04 SoftToO HD93521 ALHd120140 - ALHd120145 Was only able to observe the standard star due to the calima causing a bad seeing at the end of the night. During the standard star the seeing was around 1.3 arcsec Observed
68-505 2024-04-13 01:02 SoftToO HD93521, SN2024cld, 2024gep, 2022prr
(failed)
ALHd130006 - ALHd130036, ALHd130041 - ALHd130044 Every target was attempted, and most of them were observed successfully. All the spectroscopy observations were successful, but only one of the photometry OBs (SN2024cld) met the seeing requirements sufficiently. Spectroscopy of 2024gep was tried twice as first time was halted by rising wind speed. The first attempt hasn't been taken account in the time spent. Observed
68-505 2024-04-14 00:11 SoftToO 2024gep, SN2024cld ALHd140080 - ALHd140082, ALHd140089 - ALHd140089 The observations technically failed because seeing requirements weren't met. I'd still have a look at the data, at least for SN2024cld, because the target in the finding chart was clearly visible in the raw image. Failed
68-505 2024-04-15 00:36 SoftToO HD93521, 2024gep ALHd150103 - ALHd150118 The targets were observed successfully in a row during 2'' seeing. Observed
68-505 2024-04-19 01:43 SoftToO 2019lqo, 2024gep, BD332642 ALHd190156 - ALHd190165, ALHd190231 - ALHd190251 Observed
68-505 2024-04-20 01:00 SoftToO SN2024cld, BD332642, 2024gep, AT2022fpx ALHd200071 - ALHd200082, ALHd200087 - ALHd200092, ALHd200097 - ALHd200103 Most of the targets were observed with thin clouds and variable seeing between 1.3-2.0''. AT2022fpx was observed with clear skies and subarcsec seeing. Observed
68-505 2024-04-26 01:30 SoftToO SN2024gzk, SN2024cld, BD332642 ALHd260296 - ALHd260306, ALHd260350 - ALHd260359, ALHd260376 - ALHd260385 Observed in clear skies with seeing around 0.8" Observed
68-505 2024-04-30 00:17 SoftToO BD332642, 2024gep ALHd300096 - ALHd300100, ALHd300106 - ALHd300116 Standard star and photometry were successfull, clear sky around 1 arcsec seeing. Seeing got worse during spectroscopy which is why it is set failed. Observed
68-505 2024-05-02 01:00 SoftToO BD332642, 2024gep, SN2024cld ALHe020187 - ALHe020191, ALHe020214 - ALHe0202124, ALHe020234 - ALHe020243 Observed in clear skies or thin clouds, seeing around 1.0" for the high target, and 2.0" for the low target. As requested, the standard was observed in both 1.0" and 1.3" slits, as both were used to observe the targets Observed
68-505 2024-05-03 00:52 SoftToO SN2024gzk, BD332642 ALHe030096 - ALHe030111 Observed in clear skies with seeing below 1.0" Observed
68-505 2024-05-04 00:09 SoftToO SN2024cld ALHe040094 - ALHe040098 Observed in clear skies with seeing around 0.7" Observed
68-505 2024-05-05 00:26 SoftToO SN2024cld, BD332642 ALHe050126 - ALHe050130, ALHe050135 - ALHe050139 Clear skies, seeing of around 0.5 arcsec. The fasttrack OB from 69-406 was observed as these targets were observed. Observed
68-505 2024-05-09 00:39 SoftToO SN2024cld, BD332642 ALHe090268 - ALHe090277, ALHe090294 - ALHe090298 All targets observed with thin clouds or clearer with seeing around 1 arcsec. Observed
68-505 2024-05-13 00:42 SoftToO SN2024gzk ALHe130072 - ALHe130076, ALHe130078 - ALHe130082 Target observed with clear skies, roughly 1.5" - 2" seeing. Observed
68-505 2024-05-14 00:53 SoftToO SN2024cld, SN2024hpj ALHe140093 - ALHe140102,ALHe140112 - ALHe140116 Observed with thin clouds in < 1.3" seeing. Shared 2024hpj will be shared with proposal 501. Observed
68-505 2024-05-19 00:15 SoftToO SN2024cld, BD332642 ALHe190101 - ALHe190110 Observed with clear skies and sub-arcsec seeing. Observed
68-505 2024-05-20 02:39 SoftToO SN2024gzk, SN2024cld, SN2024hpj,
BD332642
ALHe200063 - ALHe200073, ALHe200083 - ALHe200089, ALHe200107 - ALHe200126 Thin clouds, seeing of less than 1 arcsec. Had a lot of trouble with the clouds at one point, hence the amount of targets that was CTRL-C'ed. But observations were successful in the end. Observed
68-505 2024-05-21 00:32 SoftToO SN2023aew ALHe210071 - ALHe210090 Observed with clear sky and seeing around or bit above 1 arcsec. First attempt interrupted due to clouds, but few first images might be useful. Observed
68-505 2024-05-27 01:32 SoftToO Object, SN2024cld, SP1550+330, SN2024hpj ALHe270099 - ALHe270103, ALHe270140 - ALHe270156, ALHe270177 - ALHe270183 Observed in clear skies with seeing below 1.0" Observed
68-505 2024-05-28 01:49 SoftToO SN2024gzk, BD332642, SN2024hpj ALHe280151 - ALHe280161, ALHe280229 - ALHe280233, ALHe280301 - ALHe280305 Observed in clear skies with seeing around or below 1.0" Observed
68-505 2024-06-03 00:33 SoftToO 2023ldh ALHf030192 - ALHf030204 Clear skies and a seeing of less than 1 arcsec. Unfortunately due to a rather filled night only one target was able to be observed. Observed
68-505 2024-06-07 01:03 SoftToO SN2024cld, SN2024hpj ALHf070158 - ALHf070162, ALHf070163 - ALHf070167, ALHf070187 - ALHf070193 Unable to observe all requested SoftToO's this night due to a tight schedule. A few targets were managed with clear skies and seeing ~1 arcsec. Observed
68-505 2024-06-08 00:09 SoftToO SN2024gzk ALHf080155 - ALHf080159 SN2024gzk observed succesfully in clear skies and sub arcsec seeing. Observed
68-505 2024-06-09 00:15 SoftToO AT2022fpx ALHf090266 - ALHf090271 Could only observe one target due to thick clouds covering the sky for the first part of the night. The observed target was observed with clear skies and seeing between 1 and 1.3 arcsec. Observed
68-505 2024-06-15 02:27 SoftToO SN2024kce, SN2024cld, SN2024hpj ALHf150050 - ALHf150056, ALHf150057 - ALHf150063, ALHf150073 - ALHf150087, ALHf150110 - ALHf150116 All science targets observed. Please use standard from 68-501, it is identical to the one you requested. Observed
68-505 2024-06-17 00:14 SoftToO SN2024kce ALHf170169 - ALHf170189 I suspect that the finding chart was not from the same field as the coordinates in the OB. However, the field in the Visplot tool, Aladin, and the image matched. So, the telescope was pointing at the requested coordinates. Sub-arcsecond seeing. Observed
68-505 2024-06-20 01:36 SoftToO SN2024gzk, SN2024kce, BD332642 ALHf200151 - ALHf200161, ALHf200165 - ALHf200169, ALHf200182 - ALHf200186 Observed with clear skies and sub-arcsec seeing. Observed
68-505 2024-06-21 01:18 SoftToO BD332642, SN2024hpj ALHf210107 - ALHf210111, ALHf210145 - ALHf210156 The observations were completed as requested, although I did not see the target in u' frames. I remember seeing an H-alpha line in the spectrum. Observed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: April 01 2024